OpenStudy (anonymous):

(x) = x 2 + 1 for f(2).

OpenStudy (anonymous):

hi! welcome to openstudy! :D

OpenStudy (anonymous):

so to solve, you would plug 2 in for x. f(x) = (2)^2 +1 f(x) = 4 + 1 f(x) = 5 Thus, your solution is 5. hope this helps. :D